OpenStudy (leila):

Rewrite as a sum or difference sin(3t)sin(5t)

OpenStudy (leila):

Anyonneee?

OpenStudy (gw2011):

I'm not sure what exact answer you are looking for, but here is something that involves a difference: Let A=3t Let B=5t Since cos(A-B)=(cosA)(cosB)+(sinA)(sinB) Then (sinA)(sinB)=cos(A-B)-(cosA)(cosB) and sin(3t)sin(5t)=cos(3t-5t)-cos(3t)cos(5t)